OpenStudy (anonymous):

Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

mathslover (mathslover):

2x and 3x are making diff. here .

OpenStudy (anonymous):

x+1 is a factor for both the Q.Es.... x2+3x+2=(x+1)(x+2) x2+2x+1=(x+1)^2

OpenStudy (anonymous):

Thanks!(: